OpenStudy (anonymous):

evaluate the following expression for a = 4 and b = -5 a^2 -10/2 + b

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

16-5-5=16-10=??????

OpenStudy (anonymous):

oh 6 lol

OpenStudy (alexandervonhumboldt2):

eayh